OpenStudy (anonymous):

I could really use some help understanding how to do this.. please and thanks! Model the function rule y = -2x - 3 with a table of values

jimthompson5910 (jim_thompson5910):

Plug in x = 0 y = -2x - 3 y = -2(0) - 3 y = 0 - 3 y = -3 So when x = 0, y = -3 ------------------------------ Plug in x = 1 y = -2x - 3 y = -2(1) - 3 y = -2 - 3 y = -5 So when x = 1, y = -5 ------------------------------ Plug in x = 2 y = -2x - 3 y = -2(2) - 3 y = -4 - 3 y = -7 So when x = 2, y = -7 ------------------------------ Plug in x = 3 y = -2x - 3 y = -2(3) - 3 y = -6 - 3 y = -9 So when x = 3, y = -9 ------------------------------ Plug in x = 4 y = -2x - 3 y = -2(4) - 3 y = -8 - 3 y = -11 So when x = 4, y = -11 ------------------------------ Plug in x = 5 y = -2x - 3 y = -2(5) - 3 y = -10 - 3 y = -13 So when x = 5, y = -13 ------------------------------

jimthompson5910 (jim_thompson5910):

So we get the table of values x y ----------------------- 0 -3 1 -5 2 -7 3 -9 4 -11 5 -13

OpenStudy (anonymous):

Thank you very much, that was VERY helpful! Now I know what I'm doing :)

jimthompson5910 (jim_thompson5910):

that's great, I'm very glad I could be of help to you.